# Twisted Spectra Revisited

### v1

@@ Introduction
The functor above sends a map $(A,\tau)\to(B,\sigma)$ in $\mathcal{S}_{/B\operatorname{Pic}(\mathbb{S})}$ to the functor $f_{!}:\mathrm{TwSp}(A,\tau)\to\mathrm{TwSp}(B,\sigma)$. Its right adjoint $f_{*}$ is itself colimit-preserving and thus further admits a right adjoint $f_{*}$. In particular, since the functor is symmetric monoidal, we have a natural equivalence
@@
of $\infty$-categories. Here, $\operatorname{Th}(\Omega\sigma)$ denotes the Thom spectrum of the map $\Omega\sigma:\Omega A\to\operatorname{Pic}(\mathbb{S})$.
@@ Acknowledgements
We would like to thank Ben Antieau, Stefan Behrens, Bastiaan Cnossen, Chris Douglas, David Gepner, Peter Haine, Rune Haugseng, Thomas Kragh, Achim Krause, Ciprian Manolescu, Thomas Nikolaus, Maxime Ramzi, Bertrand Toën, and Lior Yanovski, for various comments, questions, and help during the writing of this paper. Special thanks are due to Maxime Ramzi for the idea of the proof of Proposition  3.7.4. TM would like to acknowledge the support of the Giorgio and Elena Petronio Fellowship at the Institute for Advanced Study.
@@ 1.5. Parametrized categories and ambidexterity
Again in this case, the inclusion $\Pr^{R}_{\mathrm{St}}\to\operatorname{Cat}_{\infty}$ preserves all small limits [HTT]*Theorem 5.5.3.18, so in order to compute the colimit of $F$, we find ourselves computing the limit of the diagram $B^{\mathrm{op}}\to\Pr^{R}_{\mathrm{St}}\to\operatorname{Cat}_{\infty}$. Since $B$ is a Kan complex, there is an equivalence $\alpha:B\simeq B^{\mathrm{op}}$, identifying each path $\iota^{\mathrm{op}}:x_{1}\to x_{0}$ in $B^{\mathrm{op}}$ with $\iota^{-1}:x_{1}\to x_{0}$ in $B$. Thus for example, one has equivalences $\iota_{*}\simeq(\iota^{-1})^{*}$ and $\iota^{*}\simeq{\iota_{-1}}_{*}$, so that $F:B\to\Pr^{L}_{\mathrm{St}}$ factors through the subcategory $\Pr^{L,R}_{\mathrm{St}}$ consisting of functors which are simultaneously left and right adjoints. The same can be said for for $F^{\mathrm{op}}$. As a result, we are able to identify the diagrams $F$ and $F^{\mathrm{op}}$ in $\operatorname{Cat}_{\infty}$. Putting all this together, we obtain the equivalence as wanted. ∎
@@ 2.3. Haunts as categorified line bundles
We remark that $B\operatorname{Pic}(\mathbb{S})$ may itself be written as the geometric realization of the colimit diagram $B\operatorname{Pic}_{\bullet}$ which is nerve of the map *$*\to B\operatorname{Pic}(\mathbb{S})$. Thus, the map $B\to B\operatorname{Pic}(\mathbb{S})$ is equivalent to a map of colimit diagrams $N_{\bullet}(p)\to B\operatorname{Pic}_{\bullet}$.
@@ 2.4. Examples of haunts
In this section, we will construct some explicit examples of haunts. Recall that in order to construct examples of haunts on a space $B$, it is often enough to give glueing data on intersections $U_{i}\cap U_{j}$ of a trivializing open cover $\{U_{i}\}_{i\in I}$ in the form of a parametrized invertible spectrum.
@@
Here, the notation in the second factor just refers to the ordinary action the fundamental group on the universal cover and the notation $a(\gamma)$ in the third factor comes from remembering that $a\in H^{1}(Y;\mathbb{R})/H^{1}(Y;\mathbb{Z})$ and that $H^{1}(Y;\mathbb{R})\cong\operatorname{Hom}(\pi_{1}(Y),\mathbb{R})$. We remark that the Poincaré bundle should be viewed as a family of flat line bundles over $Y$ parametrized by $P$. Indeed, the following statements can be used to characterise it:
@@
With this bundle in the back of our head, let us construct the wanted K-theory class. The data of the complex spin structure $\mathfrak{s}$ can alternatively we phrased in terms of a hermitian rank $2$ bundle $S\to Y$, the spinor bundle, which comes equipped with Clifford multiplication. Note that we can consider the isomorphism class $[S]$ of the spinor bundle as an object of $K^{0}(Y)$ and that we can consider the isomorphism class $[\mathcal{L}]$ of the Poincaré bundle as an object of $K^{0}(P\times Y)$. Let us consider the projections $p:P\times Y\longrightarrow P$ and $q:P\times Y\longrightarrow Y$. Note that we have a pullback map
@@ 3.1. Twisted Spectra
in $\operatorname{Fun}(U_{i}\cap U_{j},\operatorname{Sp})$, where $\Phi_{ij}$ is an autoequivalence of $\operatorname{Fun}(U_{i}\cap U_{j},\operatorname{Sp})$. Furthermore, these autoequivalences are determined by invertible spectra satisfying a cocycle condition, so that in fact
@@ 3.2. Examples of Twisted Spectra
[…]
@@ 3.3. Pullbacks along maps of spaces
[…]
@@ 3.5. Projection isomorphism
[…]
@@ 3.7. Thom Spectrum Perspective
[…]
@@
[…]
@@
[…]
@@
[…]
@@
Natural transformations $\mathcal{C}_{/x}\to f(x)$ can be identified with (non-Cartesian) sections of the cocartesian fibration $F$ classifying $f$. In fact we see that in the $\infty$-category of sections $\Gamma(\mathcal{C},F)$, terminal objects may be understood pointwise, and so the section sending each $x$ to the terminal object of $f(x)$ is itself terminal in this category, hence determined uniquely up to contractible choice. ∎
@@
[…]
@@
[…]
@@ 4.1. Non-existence of global homotopy invariants
In particular, the equivalence $0\simeq p^{*}(0)\simeq\tau$ is forced upon us. This many not appear so interesting at first glance. Something slightly interesting is hidden in the notation though. Indeed, while the above only make sense when the twist is nullhomotopic, the invariants that we get depend on how we trivialise the twist. Note that the trivialisation corresponds to a particular equivalence $\mathrm{TwSp}(A,\tau)\simeq\operatorname{Sp}^{A}$ with the category of spectra parametrized over $A$.
@@ 4.2. Homology and Cohomology Invariants
Now, it could of course happen that we are dealing with a twisted spectrum with a non-trivializable twisting $\tau$, in which case our twisted spectrum does not have a globally defined homotopy invariant. We are not completely out of luck, though. Indeed, we might still have (co)homology invariants with respect to other generalized cohomology theories. Let $R$ be an $\mathbb{E}_{\infty}$-ring, so that we have an induced map $B\operatorname{Pic}(\mathbb{S})\to B\operatorname{Pic}(R)$ of Picard spaces. Now, if the composition $\tau_{R}:A\to B\operatorname{Pic}(\mathbb{S})\to B\operatorname{Pic}(R)$ is nullhomotopic, then it makes sense to talk about $R$-homology and $R$-cohomology of $X$, completely analogous to before. Again the (co)homology will depend on our choice of trivialisation of $\tau_{R}$, though.
@@
A haunt $\tau:P\to B\operatorname{Pic}(\mathbb{S})$ obtained from a complex spin manifold $(Y,\mathfrak{s})$ always admits $\operatorname{KU}$-invariants. Indeed, for dimension reasons on $P$, we only need to understand the induced map
@@
for $i=1,2,3$. It is clear that the map is zero for $i=2$ since $\pi_{2}U\cong 0$, and that it is zero for $i=3$ since $\pi_{3}B\operatorname{Pic}(\operatorname{KU})\cong 0$. We also claim that the map induced for $i=1$ is zero. The rough argument uses that isomorphism classes of complex spin structures on $Y$ is in one-to-one correspondence with complex line bundles over $Y$ [KM07]*Proposition 1.1.1. Indeed, given one complex spin structure on $Y$ with spinor bundle $S$, all other complex spin structures can be obtained by tensoring $S\otimes L$ with a hermitian line bundle over $Y$. The first Chern character of $S\otimes L$ is known to be
@@ 5.2. The Total Pullback Functors and Their Adjoints
Note that we can think of all the pullback functors functors $f^{*}:\mathrm{TwSp}(B,\tau)\to\mathrm{TwSp}(A,f^{*}\tau)$ for various twists $\tau$ as assembling into a natural transformation
@@ 5.4. Projection Isomorphism on Total Category
[…]
@@ 5.5. Closed Structure on the Total Category
[…]
@@ 6.1. The $(\infty,2)$-category of twisted parametrized spectra
$\operatorname{\mathbf{TwSp}}$
$\operatorname{\mathbf{Pr}}_{\mathrm{St}}^{L}$
objects
twist $\tau:A\to B\operatorname{Pic}(\mathbb{S})$
presentable stable $\infty$-category
@@
1-morphism
$X\in\mathrm{TwSp}(A\times B,-\tau\boxtimes\sigma)$
colimit preserving functor
@@
2-morphism
map $X\to Y$ of twisted spectra
natural transformation
@@
identity 1-morphism
$\tau$-twisted spherical group ring
@@
vertical composition
composition of maps of twisted spectra
vertical composition of natural transformations
horizontal composition of 1-morphisms
$Y\odot X\in\mathrm{TwSp}(A\times C,-\tau\boxtimes\rho)$
composition of functors
@@
[…]
@@
$\Phi_{Y\odot X}$ $*\displaystyle=(r_{C})_{!}\circ(-\otimes(Y\odot X))\circ(r_{A})^{*}$
$*.\displaystyle=(r_{C})_{!}\circ(-\otimes((p_{A\times C})_{!}(p^{*}_{B\times C}Y\otimes p_{A\times B}^{*}X)))\circ(r_{A})^{*}\,.$
@@
[…]
@@
$*.\Phi_{Y}\circ\Phi_{X}\simeq(q_{C})_{!}\circ(-\otimes Y)\circ(p_{B\times C})_{!}\circ(-\otimes(p_{A\times B})^{*}(X))\circ(\pi_{A})^{*}\,.$
@@
$*.\Phi_{Y}\circ\Phi_{X}\simeq(\pi_{C})_{!}\circ(-\otimes(p_{B\times C})^{*}(Y))\circ(-\otimes(p_{A\times B})^{*}(X))\circ(\pi_{A})^{*}\,.$
@@
$*\Phi_{Y\odot X}\simeq(\pi_{C})_{!}\circ(-\otimes((p_{B\times C})^{*}Y\otimes(p_{A\times B})^{*}X))\circ(\pi_{A})^{*}$
@@
$*\displaystyle(p_{2})_{!}\circ(-\otimes\Delta_{!}\mathbb{S})\circ(p_{1})^{*}$ $*\displaystyle\simeq(p_{2})_{!}\circ\Delta_{!}\circ(-\otimes\mathbb{S})\circ\Delta^{*}(p_{1})^{*}$
@@
$\Phi_{(1,f)_{!}\mathbb{S}}$ $*\displaystyle=(p_{B})_{!}\circ(-\otimes(1,f)_{!}\mathbb{S})\circ(p_{A})^{*}$
$*\displaystyle\simeq(p_{B})_{!}\circ(1,f)_{!}\circ(-\otimes\mathbb{S})\circ(1,f)^{*}\circ(p_{A})^{*}$

### v2

@@ Introduction
The functor above sends a map $f:(A,\tau)\to(B,\sigma)$ in $\mathcal{S}_{/B\operatorname{Pic}(\mathbb{S})}$ to the functor $f_{!}:\mathrm{TwSp}(A,\tau)\to\mathrm{TwSp}(B,\sigma)$. Its right adjoint $f^{*}$ is itself colimit-preserving and thus further admits a right adjoint $f_{*}$. In particular, since the functor is symmetric monoidal, we have a natural equivalence
@@
of $\infty$-categories. Here, $\operatorname{Th}(\Omega\sigma)$ denotes the Thom spectrum of the map $\Omega\sigma:\Omega B\to\operatorname{Pic}(\mathbb{S})$.
@@ Acknowledgements
We would like to thank Ben Antieau, Stefan Behrens, Bastiaan Cnossen, Chris Douglas, David Gepner, Peter Haine, Rune Haugseng, Thomas Kragh, Achim Krause, Ciprian Manolescu, Thomas Nikolaus, Maxime Ramzi, Bertrand Toën, and Lior Yanovski, for various comments, questions, and help during the writing of this paper. Special thanks are due to Maxime Ramzi for the idea of the proof of Proposition  3.7.4. TM would like to acknowledge the support of the Giorgio and Elena Petronio Fellowship at the Institute for Advanced Study. We would also like to thank an anonymous referee for helpful comments on an earlier version of this document.
@@ 1.5. Parametrized categories and ambidexterity
Again in this case, the inclusion $\Pr^{R}_{\mathrm{St}}\to\operatorname{Cat}_{\infty}$ preserves all small limits [HTT]*Theorem 5.5.3.18, so in order to compute the colimit of $F$, we find ourselves computing the limit of the diagram $B^{\mathrm{op}}\to\Pr^{R}_{\mathrm{St}}\to\operatorname{Cat}_{\infty}$. Since $B$ is a Kan complex, there is an equivalence $\alpha:B\simeq B^{\mathrm{op}}$, identifying each path $\iota^{\mathrm{op}}:x_{1}\to x_{0}$ in $B^{\mathrm{op}}$ with $\iota^{-1}:x_{1}\to x_{0}$ in $B$. Since the functor $F$ factors through the core groupoid of $\mathrm{Pr}^{L}_{\mathrm{St}}$, and the natural diagram
[…]
commutes, we are able to identify the diagrams $F$ and $F^{\mathrm{op}}$ in $\operatorname{Cat}_{\infty}$. Putting all this together, we obtain the equivalence as wanted. ∎
@@ 1.6. Locally constant sheaves and parametrized objects
The above theorem holds true even if we replace $\operatorname{Cat}_{R,\kappa}$ with any $\infty$-category $\mathcal{C}$ admitting small limits, with a different proof.
Remark 1.6.4.
@@ 2.3. Haunts as categorified line bundles
We remark that $B\operatorname{Pic}(\mathbb{S})$ may itself be written as the geometric realization of the colimit diagram $B\operatorname{Pic}_{\bullet}$ which is nerve of the map $*\to B\operatorname{Pic}(\mathbb{S})$. Thus, the map $B\to B\operatorname{Pic}(\mathbb{S})$ is equivalent to a map of colimit diagrams $N_{\bullet}(p)\to B\operatorname{Pic}_{\bullet}$.
@@ 2.4. Examples of haunts
In this section, we will construct some explicit examples of haunts. Recall that in order to construct examples of haunts on a space $B$, it is often enough to give gluing data on intersections $U_{i}\cap U_{j}$ of a trivializing open cover $\{U_{i}\}_{i\in I}$ in the form of a parametrized invertible spectrum.
@@
Here, the notation in the second factor just refers to the ordinary action the fundamental group on the universal cover and the notation $a(\gamma)$ in the third factor comes from remembering that $a\in H^{1}(Y;\mathbb{R})/H^{1}(Y;\mathbb{Z})$ and that $H^{1}(Y;\mathbb{R})\cong\operatorname{Hom}(\pi_{1}(Y),\mathbb{R})$. We remark that the Poincaré bundle should be viewed as a family of flat line bundles over $Y$ parametrized by $P$. Indeed, the following statements can be used to characterize it:
@@
With this bundle in the back of our head, let us construct the wanted K-theory class. The data of the complex spin structure $\mathfrak{s}$ can alternatively be phrased in terms of a hermitian rank $2$ bundle $S\to Y$, the spinor bundle, which comes equipped with Clifford multiplication. Note that we can consider the isomorphism class $[S]$ of the spinor bundle as an object of $K^{0}(Y)$ and that we can consider the isomorphism class $[\mathcal{L}]$ of the Poincaré bundle as an object of $K^{0}(P\times Y)$. Let us consider the projections $p:P\times Y\longrightarrow P$ and $q:P\times Y\longrightarrow Y$. Note that we have a pullback map
@@ 3.1. Twisted Spectra
in $\operatorname{Fun}(U_{i}\cap U_{j},\operatorname{Sp})$, where $\Phi_{ij}$ is an auto-equivalence of $\operatorname{Fun}(U_{i}\cap U_{j},\operatorname{Sp})$. Furthermore, these auto-equivalences are determined by invertible spectra satisfying a cocycle condition, so that in fact
@@ 3.2. Examples of Twisted Spectra
${\mathrm{TwSp}(S^{1},n)}$${\operatorname{Sp}}$${\operatorname{Sp}}$${\operatorname{Sp}\times\operatorname{Sp}}$
@@ 3.3. Pullbacks along maps of spaces
$f_{!}:\mathrm{TwSp}(A,f^{*}\sigma)\longrightarrow\mathrm{TwSp}(B,\sigma)\quad\text{and}\quad f_{*}:\mathrm{TwSp}(A,f^{*}\sigma)\longrightarrow\mathrm{TwSp}(B,\sigma)\,,$
@@ 3.5. Projection isomorphism
${(A,f^{*}\sigma+f^{*}\rho)}$${(B\times A,\sigma\boxtimes f^{*}\rho)}$${(B,\sigma+\rho)}$${(B\times B,\sigma\boxtimes\rho)\,,}$$\scriptstyle{f}$$\scriptstyle{(f,1)}$$\scriptstyle{1\times f}$$\scriptstyle{\Delta_{B}}$
@@ 3.7. Thom Spectrum Perspective
${\operatorname{Sp}}$${\mathrm{TwSp}(A,f^{*}\sigma)}$${\mathrm{TwSp}(B,\sigma)}$$\scriptstyle{a_{!}}$$\scriptstyle{b_{!}}$$\scriptstyle{f_{!}}$
@@
${\operatorname{Mod}_{\operatorname{End}(a_{!}\mathbb{S})}}$${\mathrm{TwSp}(A,f^{*}\sigma)}$${\operatorname{Mod}_{\operatorname{End}(b_{!}\mathbb{S})}}$${\mathrm{TwSp}(B,\sigma)\,.}$$\scriptstyle{\operatorname{End}(f_{!}\mathbb{S})_{!}}$$\scriptstyle{f_{!}}$
@@
$\operatorname{Th}:\operatorname{Grp}_{/\operatorname{Pic}(\mathbb{S})}\longrightarrow\operatorname{Alg}\quad\text{and}\quad\operatorname{End}((B\iota_{(-)})_{!}\mathbb{S}):\operatorname{Grp}_{/\operatorname{Pic}(\mathbb{S})}\longrightarrow\operatorname{Alg}\,.$
@@
(4) ${\operatorname{Grp}_{/\operatorname{Pic}(\mathbb{S})}}$${(\mathrm{Pr}^{L}_{\mathrm{St}})_{\operatorname{Sp}/}}$${\operatorname{Alg}}$${\operatorname{Mod}_{\mathcal{S}[\operatorname{Pic}(\mathbb{S})]}(\mathrm{Pr}^{L})_{\mathcal{S}[\operatorname{Pic}(\mathbb{S})]/}}$${\operatorname{Alg}(\mathcal{S}[\operatorname{Pic}(\mathbb{S})])\,.}$$\scriptstyle{\operatorname{End}(1_{(-)})}$$\scriptstyle{\operatorname{End}^{\mathcal{S}[\operatorname{Pic}(\mathbb{S})]}(1_{(-)})}$$\scriptstyle{(\epsilon_{\operatorname{Sp}})_{!}}$$\scriptstyle{\operatorname{Alg}(\epsilon_{\operatorname{Sp}})}$
@@
Natural transformations $\mathcal{C}_{/(-)}\to f$ can be identified with (non-Cartesian) sections of the cocartesian fibration $F$ classifying $f$. Informally, the section $\sigma$ sends some element $y$ to a pair $(y,\beta(y)\in f(y))$, from which the natural transformation $\tau:\mathcal{C}_{/(-)}\to f$ corresponding to $\sigma$ is such that
$\tau_{x}:\mathcal{C}_{/x}\to f(x)\,,\quad(g:y\to x)\mapsto g_{*}(\beta(y))\,,$
where $g_{*}:f(y)\to f(x)$ is the functor induced by evaluating $f$ on $g$. Since we are assuming that we are dealing with natural transformations that send the identity to terminal objects, we must then have that $\beta(y)$ is terminal in $f(y)$ for every $y$. Since terminal objects in the $\infty$-category of sections $\Gamma(\mathcal{C},F)$ may be understood pointwise, it follows that the section $\sigma$ itself is terminal in this $\infty$-category. Hence it is determined uniquely up to contractible choice. ∎
@@
${\Omega A}$${\Omega B}$${\operatorname{Pic}(\mathbb{S})}$$\scriptstyle{\Omega f}$$\scriptstyle{\Omega(f^{*}\sigma)}$$\scriptstyle{\Omega\sigma}$
@@
${\operatorname{Th}(\Omega(f^{*}\sigma))}$${\operatorname{Th}(\Omega\sigma)}$${\operatorname{End}(a_{!}\mathbb{S})}$${\operatorname{End}(b_{!}\mathbb{S})}$$\scriptstyle{\operatorname{Th}(\Omega f)}$$\scriptstyle{\operatorname{End}(f_{!}\mathbb{S})}$
@@ 4.1. Non-existence of global homotopy invariants
In particular, the equivalence $0\simeq p^{*}(0)\simeq\tau$ is forced upon us. This many not appear so interesting at first glance. Something slightly interesting is hidden in the notation though. Indeed, while the above only make sense when the twist is nullhomotopic, the invariants that we get depend on how we trivialize the twist. Note that the trivialization corresponds to a particular equivalence $\mathrm{TwSp}(A,\tau)\simeq\operatorname{Sp}^{A}$ with the category of spectra parametrized over $A$.
@@ 4.2. Homology and Cohomology Invariants
Now, it could of course happen that we are dealing with a twisted spectrum with a non-trivializable twisting $\tau$, in which case our twisted spectrum does not have a globally defined homotopy invariant. We are not completely out of luck, though. Indeed, we might still have (co)homology invariants with respect to other generalized cohomology theories. Let $R$ be an $\mathbb{E}_{\infty}$-ring, so that we have an induced map $B\operatorname{Pic}(\mathbb{S})\to B\operatorname{Pic}(R)$ of Picard spaces. Now, if the composition $\tau_{R}:A\to B\operatorname{Pic}(\mathbb{S})\to B\operatorname{Pic}(R)$ is nullhomotopic, then it makes sense to talk about $R$-homology and $R$-cohomology of $X$, completely analogous to before. Again the (co)homology will depend on our choice of trivialization of $\tau_{R}$, though.
@@
A haunt $\tau:P\to B\operatorname{Pic}(\mathbb{S})$ obtained from a complex spin manifold $(Y,\mathfrak{s})$ always admits $\operatorname{KU}$-invariants. Note that the induced map
@@
is zero for all $i$. Indeed, it is clear that the map is zero for even $i\geq 2$ since $\pi_{i}U\cong 0$ for all even $i$, and that it is zero for odd $i\geq 3$ since $\pi_{i}B\operatorname{Pic}(\operatorname{KU})\cong 0$ for all odd $i$ in this range. We also claim that the map induced for $i=1$ is zero. The rough argument uses that isomorphism classes of complex spin structures on $Y$ is in one-to-one correspondence with complex line bundles over $Y$ [KM07]*Proposition 1.1.1. Indeed, given one complex spin structure on $Y$ with spinor bundle $S$, all other complex spin structures can be obtained by tensoring $S\otimes L$ with a hermitian line bundle over $Y$. The first Chern character of $S\otimes L$ is known to be
@@
Moreover, we want to point out the following result of Maegawa which tells us that twisted spectra coming from haunts that factor through the complex J-homomorphism, which are common in Floer homotopy theory, always admit $R$-(co)homology group with respect to any even $\mathbb{E}_{2}$-ring $R$.
Proposition 4.2.6 ([maegawa24]*Theorem 3.1).
Let $B$ be a CW-complex and let $\tau:B\to B(\mathbb{Z}\times\operatorname{BU})\to B\operatorname{Pic}(\mathbb{S})$ be a haunt that factor through the complex J-homomorphism. Then $\tau_{R}:B\to B\operatorname{Pic}(R)$ is nullhomotopic for any even periodic $\mathbb{E}_{2}$-ring spectrum. Moreover, the nullhomotopy is determined by essentially uniquely by the choice of an $\mathbb{E}_{1}$-complex orientation on $R$.
@@ 5.2. The Total Pullback Functors and Their Adjoints
Note that we can think of all the pullback functors $f^{*}:\mathrm{TwSp}(B,\tau)\to\mathrm{TwSp}(A,f^{*}\tau)$ for various twists $\tau$ as assembling into a natural transformation
@@ 5.4. Projection Isomorphism on Total Category
${A}$${B\times A}$${B}$${B\times B}$$\scriptstyle{f}$$\scriptstyle{(f,1)}$$\scriptstyle{1\times f}$$\scriptstyle{\Delta_{B}}$
@@ 5.5. Closed Structure on the Total Category
${\mathrm{TwSp}(A)}$${\mathrm{TwSp}(A)}$${\mathrm{TwSp}(A,,\tau+\sigma)}$${\mathrm{TwSp}(A,\tau)\,,}$$\scriptstyle{\hom(X,-)}$$\scriptstyle{\hom(X,-)}$
@@ 6.1. The $(\infty,2)$-category of twisted parametrized spectra
$\operatorname{\mathbf{TwSp}}$ $\operatorname{\mathbf{Pr}}_{\mathrm{St}}^{L}$
objects twist $\tau:A\to B\operatorname{Pic}(\mathbb{S})$ presentable stable $\infty$-category
@@
1-morphism $X\in\mathrm{TwSp}(A\times B,-\tau\boxtimes\sigma)$ colimit preserving functor
@@
2-morphism map $X\to Y$ of twisted spectra natural transformation
@@
identity 1-morphism $\tau$-twisted spherical group ring
@@
vertical composition composition of maps of twisted spectra vertical composition of natural transformations
horizontal composition of 1-morphisms $Y\odot X\in\mathrm{TwSp}(A\times C,-\tau\boxtimes\rho)$ composition of functors
@@
$\Phi_{Y\odot X}\simeq\Phi_{Y}\circ\Phi_{X}\quad\text{and}\quad\Phi_{\Delta^{\tau}_{!}\mathbb{S}}\simeq 1_{\mathrm{TwSp}(A,\tau)}\,.$
@@
$\Phi_{Y\odot X}$ $=(r_{C})_{!}\circ(-\otimes(Y\odot X))\circ(r_{A})^{*}$
$=(r_{C})_{!}\circ(-\otimes((p_{A\times C})_{!}(p^{*}_{B\times C}Y\otimes p_{A\times B}^{*}X)))\circ(r_{A})^{*}\,.$
@@
${A\times B\times C}$${A\times B}$${B\times C}$${B}$$\scriptstyle{p_{A\times B}}$$\scriptstyle{p_{B\times C}}$$\scriptstyle{p_{B}}$$\scriptstyle{q_{B}}$
@@
$\Phi_{Y}\circ\Phi_{X}\simeq(q_{C})_{!}\circ(-\otimes Y)\circ(p_{B\times C})_{!}\circ(-\otimes(p_{A\times B})^{*}(X))\circ(\pi_{A})^{*}\,.$
@@
$\Phi_{Y}\circ\Phi_{X}\simeq(\pi_{C})_{!}\circ(-\otimes(p_{B\times C})^{*}(Y))\circ(-\otimes(p_{A\times B})^{*}(X))\circ(\pi_{A})^{*}\,.$
@@
$\Phi_{Y\odot X}\simeq(\pi_{C})_{!}\circ(-\otimes((p_{B\times C})^{*}Y\otimes(p_{A\times B})^{*}X))\circ(\pi_{A})^{*}$
@@
$(p_{2})_{!}\circ(-\otimes\Delta_{!}\mathbb{S})\circ(p_{1})^{*}$ $\simeq(p_{2})_{!}\circ\Delta_{!}\circ(-\otimes\mathbb{S})\circ\Delta^{*}(p_{1})^{*}$
@@
$\Phi_{(1,f)_{!}\mathbb{S}}$ $=(p_{B})_{!}\circ(-\otimes(1,f)_{!}\mathbb{S})\circ(p_{A})^{*}$
$\simeq(p_{B})_{!}\circ(1,f)_{!}\circ(-\otimes\mathbb{S})\circ(1,f)^{*}\circ(p_{A})^{*}$